OpenStudy (anonymous):

A total bill including tax, was $43.74 If the tax rate was 8% what was the original price?

OpenStudy (saifoo.khan):

Stop giving wrong answers man!

OpenStudy (saifoo.khan):

lol, jk.

OpenStudy (anonymous):

108/100 = 43.74/x 4374/108 = 40.50

OpenStudy (anonymous):

no names khan

OpenStudy (anonymous):

am i good now khan

OpenStudy (saifoo.khan):

lol, i was kiddin/.

OpenStudy (anonymous):

hahahahahhahahhahahahaha we love u bro and aukaii if it wasent for u I would have got it wrong lmao

OpenStudy (anonymous):

x+(0.08x)=43.74

OpenStudy (anonymous):

lol i know

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

1.08x=43.74 x=43.74/1.08 x=40.5

OpenStudy (saifoo.khan):

she is correct i guess.

OpenStudy (anonymous):

thank you nancy

OpenStudy (anonymous):

man khan you remind me of my self you fly threw them and fget them wrong slowwwwwwwwwwwwwwwww down lmaoooooooooooooooo

OpenStudy (anonymous):

see look I dont even know how to spell get. hahaha

OpenStudy (saifoo.khan):

lol, cant get the last part.

OpenStudy (anonymous):

will you help me with a couple more. you have time?

OpenStudy (saifoo.khan):

Sure, yeah. i have time today! :D

OpenStudy (anonymous):

How much interest will will roxanne have to pay if she borrows $2000.00 for 2 years at a simple interest rate of 4%?

OpenStudy (saifoo.khan):

I = PRT I = 2000x0.04x2 I = $160

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about this one? If 1500.00 grows to 1950.00 after 5 years , find the simple interest rate.

OpenStudy (saifoo.khan):

0.26

OpenStudy (saifoo.khan):

0.26%

OpenStudy (saifoo.khan):

Please hit good answer.